# Towards Calibrated Robust Fine-Tuning of Vision-Language Models

### v5

@@ 5.3 Empirical studies
Table 7: ImageNet accuracy and ECE results on different backbones. We provide summarized results on CLIP RN50 and ViT-L/14. The best and the second-best in each column are underlined. (See Table G and H for details.)
@@
Analysis on singular values. Figure 4 illustrates the last 20 singular values of the covariance matrix $\bar{I}^{T}\bar{I}$ where $\bar{I}$ is a standardized image representations over $N$ examples. Our proposed constrained contrastive loss $\mathcal{L}_{\text{MCL-con}}$ enlarges the small singular values compared to the conventional contrastive loss $\mathcal{L}_{\text{MCL}}$. This result verifies that adding the orthogonality constraint successfully reduces $1/\sigma_{\text{min}}(\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}})$, the component of the shared upper bound we derived in §3, following our intention.
@@ 6 Related Work
Robust fine-tuning for visual foundation models. Beyond the ID generalization, there are a lot of works aiming at improving the generalization of fine-tuned models on the OOD domain. Some of them leverage the strong robustness of pre-trained model through weight-average [53, 20, 46] or regularization [46, 49, 50, 36] whereas others attribute to the robustness during fine-tuning from different part of model backbone [25, 26]. Besides, Goyal et al. [13] claim that aligning the learning objective during pre-training and fine-tuning is crucial for retaining the remarkable OOD generalization capability of the pre-trained model. Although the above methods have provided insights into the extrapolation of foundation models, confidence calibration has been unaddressed, which is crucial for reliable ML applications. We investigate the calibration of fine-tuned CLIP and propose a unified fine-tuning strategy that has theoretical support to achieve superior ID and OOD calibration alongside OOD generalization for the first time.
Confidence calibration. After some early research on calibrated prediction [33, 7], lots of follow-up studies have been conducted. As a seminal work, Guo et al. [14] revealed the miscalibration problem of neural networks, then, Minderer et al. [29] and LeVine et al. [27] provided a comprehensive analysis on the calibration of modern vision models with consideration on distribution shift. To improve the calibration of predictive models, Temperature Scaling (TS) [14] and Label Smoothing (LS) [47] are two representative methods in practice. TS-based approaches learn a temperature parameter itself [14, 12] or model [54, 23] to estimate the temperature to adjust the output probability of models, whereas LS-based methods focus on producing soft labels to mitigating overconfidence issues by a fixed [47, 32], randomized [48], or model-based [62, 57] smoothing strategies. However, existing approaches do not consider distribution shift [62], assume accessibility to target domain [12, 54], assume specific type of distribution shift [51], cannot adjust confidences individually [14, 47]. In this work, we adopt EMA self-distillation as an effective input-dependent calibration method and show that the superior calibration results on in-domain samples can be transferred to other domains (without data from those domains) by pursuing the larger smallest singular value together.
7 Conclusion
@@
Limitation and Societal impact. Our research reached the scale of ViT-L. Due to resource constraints, we were unable to perform more experiments on larger models like ViT-G or ViT-H.
Societal impact. Our experiments utilize the ImageNet dataset, which is known for certain inherent biases, such as geographic or cultural biases. While our method achieves good generalization and calibration results that can positively impact numerous real-world applications, biases learned from the dataset could have negative downstream impacts.
@@ Appendix A Experimental Detail
This section supplements §5 providing detailed explanations for experiments to enhance the reproducibility.
@@ A.1 Details for numerical analysis on error bounds
We generate 1000-dimensional Gaussian random variables, where the variables have binary noisy labels ($15\%$ of random flip) for the ID train, ID test, and OOD test datasets. For the ID train set, the first 400 dimensions and the second 400 dimensions are correlated with labels, and the remaining 200 dimensions are zero-centered random noises. We build the OOD test set from the ID test set by shifting the mean of the first 400 dimensions and downscaling the second 400 dimensions in half. The remaining 200 dimensions are intact. For example, the feature $x_{2}$ in Figure A is perfectly correlated with labels across train and test in both ID and OOD environments, while the correlation between feature $x_{1}$ and labels vanish in OOD environment. We train the three-layer multi-layer perceptron networks with four different learning objectives adopting regularization terms for calibration ($\mathcal{L}_{\text{SD}}$) and for the smallest singular value ($\mathcal{L}_{\text{OC}}$) with varying regularization magnitudes (111 models in total): (i) without regularization (Baseline) (ii) with $\mathcal{L}_{\text{SD}}$, (iii) with $\mathcal{L}_{\text{OC}}$, and (iv) with $\mathcal{L}_{\text{SD}}$ and $\mathcal{L}_{\text{OC}}$. For $\mathcal{L}_{\text{OC}}$, we use an orthogonal constraint over the last weight matrix. For $\mathcal{L}_{\text{SD}}$, we adopt Born-Again-Network (BAN)-style self-distillation [10]. After training, we measured $\sigma_{min}$ and ECE on the ID test set and measured the mean squared error (MSE) and ECE on OOD test set.
@@ Appendix B Additional Evaluation Results
In addition to the comparisons of our CaRot with zero-shot, naive fine-tuning, and robust fine-tuning methods in §5.3, we provide results when applying post-hoc techniques for robustness (weight ensembling) and calibration (temperature scaling). Moreover, we compare our self-distillation-based soft label with uniform constant label smoothing.
@@ B.1 Comparing approaches
B.2 Additional results
Results with WiSE are reported in Table A and B. We observe that weight ensembling, which aims to align zero-shot models and fine-tuned models in the model weight space, boosts the overall performance. Still, CaRot shows superior results to the baselines.
Results with TS can be found in Table C. Aligning with previous observations, applying TS significantly improves ID ECE. However, since TS adjusts the logits assuming that all data instances would have a similar extent of overconfidence (or underconfidence) issue, its positive effect is limited under distribution shifts. Hence, TS does not guarantee to attain low calibration error on OOD datasets. It is noteworthy that applying TS even harms OOD ECE of CaRot (compare results with Table 3), which is already calibrated using data-dependent confidence adjustment during fine-tuning.
Results comparing with LS are reported in Table D. In accordance with its capability of improving calibration, LS successfully reduces calibration errors. We observe that our adopted approach, EMA SD, remarkably outperforms LS, especially on OOD ECE as well as in ID ECE. While LS addresses the over-confidence problem during train time by dispersing the concentrated confidence on the target label to non-target labels with a constant amount, self-distillation shows a similar behavior but in a dynamic approach, considering the diversity of data instances [62]. We interpret that reflecting the difficulty of input batch and distilling such information is crucial to achieving robust calibration.
@@ B.2 Additional results
Table A: Accuracy on ImageNet and distribution shifts using WiSE-FT [53]. We select the optimal ensembling coefficient (i.e., $\alpha$) for each method.
@@
Table B: ImageNet ECE results with WiSE-FT [53]. Along with Table A, we report the ECE results.
@@
Table C: ImageNet ECE results with temperature scaling (TS). Supplement to Table 2 and 3, we provide results applying TS. Note that TS is a post-hoc method and does not affect accuracy. The temperature is selected using IN ECE for each method.
@@ B.3 Detailed results from main paper
Coefficient terms ablation. In Table F and E, we present the detailed ablation results of hyperparameters associated with the methodologies addressed in our paper. These results supplement Table 5.
Different VLM backbones. We provide the full results of Table 7 in Table G and H.
@@
Table D: Comparison on LS and EMA SD. We compare the impact of LS and EMA SD with $\mathcal{L}_{\text{MCL-con}}$ as calibration regularization.
@@
Table E: Orthogonality constraint hyperparameter. We report the impact of the orthogonality constraint term of CaRot objective by ablating its strength coefficient $\lambda_{\text{OC}}$. We set our final value as 0.2, tuning based on ID ECE.
@@
Table F: Self-distillation term hyperparameter. We report the impact of the EMA SD term of CaRot objective by ablating its strength coefficient $\lambda_{\text{SD}}$. We set our final value as 1.5, tuning based on $\sigma_{min}$.
@@
Table G: ImageNet results on CLIP ResNet50
@@
Table H: ImageNet results on CLIP ViT-L/14
@@
Appendix C Proof
In this section, we provide a proof for Theorem 3.1.
@@ Appendix C Proof
Let $h:\mathcal{X}\rightarrow[0,1]$ be a real-valued function which has structure $h(x)=\sum_{i=1}^{d}h_{i}(x_{i})$ where $h_{i}$ is an arbitrary one-dimensional function, and $h(\cdot)$ is in a hypothesis class $\mathcal{H}$ that has pseudo dimension $\mathcal{P}dim(\mathcal{H})=d_{h}$, $\hat{\mathcal{D}}_{\text{ID}}$ be an $N$-size empirical distribution on ID domain. If $(x_{i},x_{j})$ is a bi-variate Gaussian random variable for every $i,j\in[d]$, then for any $\delta\in(0,1)$ and for all $h$, the following bounds hold with probability at least $1-\delta$:
@@
where $\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}}{:=}\mathbb{E}_{\mathcal{D}_{\text{ID}}}[\tilde{x}\tilde{x}^{T}]$ is a covariance matrix of $d$-dimensional normalized input $\tilde{x}=(\tilde{x}_{1},...,\tilde{x}_{d})$, where $\tilde{x}_{i}{:=}(x_{i}-\mathbb{E}[x_{i}])\text{Var}(x_{i})^{-1/2}$ and $\sigma_{\text{min}}(M)$ is the smallest singular value of a matrix $M\in\mathbb{R}^{d_{1}\times d_{2}}$.
@@
$\varepsilon_{\mathcal{D}_{\text{OOD}}}(h)$ $\leq\varepsilon_{\mathcal{D}_{\text{OOD}}}(h*)+\varepsilon_{\mathcal{D}_{\text{OOD}}}(h,h*)$
$=\varepsilon_{\mathcal{D}_{\text{OOD}}}(h*)+\varepsilon_{\mathcal{D}_{\text{OOD}}}(h,h*)-\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h*)+\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h*)$
$\leq\varepsilon_{\mathcal{D}_{\text{OOD}}}(h*)+|\varepsilon_{\mathcal{D}_{\text{OOD}}}(h,h*)-\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h*)|+\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h*)$
$\leq\varepsilon_{\mathcal{D}_{\text{OOD}}}(h*)+\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h*)+sd_{\mathcal{H}}(\mathcal{D}_{\text{OOD}},\mathcal{D}_{\text{ID}})$
$\leq\varepsilon_{\mathcal{D}_{\text{OOD}}}(h*)+\varepsilon_{\mathcal{D}_{\text{ID}}}(h)+\varepsilon_{\mathcal{D}_{\text{ID}}}(h*)+sd_{\mathcal{H}}(\mathcal{D}_{\text{OOD}},\mathcal{D}_{\text{ID}})$
@@
where $\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}}:=\mathbb{E}_{\mathcal{D}_{\text{ID}}}[\tilde{x}\tilde{x}^{T}]$ is a covariance matrix of the $d$-dimensional nomarlized input $\tilde{x}=(\tilde{x}_{1},...,\tilde{x}_{d})$ where $\tilde{x}_{i}:=(x_{i}-\mathbb{E}[x_{i}])\text{Var}(x_{i})^{-1/2}$ and $\sigma_{\text{min}}(M)$ is the minimum singular value of a matrix $M\in\mathbb{R}^{d_{1}\times d_{2}}$.

### v7

@@ 5.3 Further empirical studies
Table 7: ImageNet accuracy and ECE on different backbones. We provide summarized results on CLIP RN50 and ViT-L/14. The best and the second-best in each column are underlined. (See Table H and I for details.)
@@
Analysis on singular values. Figure 6 illustrates the last 20 singular values of the covariance matrix $\bar{I}^{T}\bar{I}$ where $\bar{I}$ is a standardized image representations over $N$ samples. Our proposed constrained contrastive loss $\mathcal{L}_{\text{MCL-con}}$ increases the small singular values compared to the vanilla contrastive loss $\mathcal{L}_{\text{MCL}}$. This result verifies that adding the orthogonality constraint successfully reduces $1/\sigma_{\text{min}}(\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}})$, the component of the shared upper bound we derived in §3, following our intention.
@@ 6 Related Work
Robust fine-tuning for visual foundation models. Beyond the ID generalization, there are a lot of works aiming at improving the generalization of fine-tuned models on the OOD domain. Some of them leverage the strong robustness of pre-trained model through weight-average [61, 26, 53, 44] or regularization [53, 57, 58, 42] whereas others attribute to the robustness during fine-tuning from different part of model backbone [30, 32]. Besides, Goyal et al. [17] claims that aligning the learning objective during pre-training and fine-tuning is crucial for retaining the remarkable OOD generalization capability of the pre-trained model. Although the above methods have provided insights into the extrapolation of foundation models regarding accuracy, confidence calibration has been unexplored, which is crucial for reliable ML applications. We investigate the OOD calibration of fine-tuned CLIP as well as accuracy and propose a unified fine-tuning strategy with theoretical support to achieve superior ID and OOD calibration alongside OOD generalization for the first time.
Confidence calibration. After some early research on calibrated prediction [39, 8], lots of follow-up studies have been conducted. As a seminal work, Guo et al. [18] revealed the miscalibration problem of neural networks, then, Minderer et al. [35] and LeVine et al. [33] provided a comprehensive analysis on the calibration of modern vision models with consideration on distribution shift. To improve the calibration of predictive models, Temperature Scaling (TS) [18] and Label Smoothing (LS) [55] are two representative methods in practice. TS-based approaches learn a temperature parameter itself [18, 16] or model [63, 28] to estimate the temperature to adjust the output probability of models, whereas LS-based methods focus on producing soft labels to mitigating overconfidence issues by a fixed [55, 38], randomized [56], or model-based [71, 66] smoothing strategies. However, existing approaches do not consider distribution shifts [71], assume accessibility to target domain [16, 63], assume specific type of distribution shift [59], cannot adjust confidences individually [18, 55]. In this work, we adopt EMA self-distillation as an effective input-dependent calibration method and show that the superior calibration results on in-domain samples can be transferred to other domains (without data from those domains) by pursuing the larger smallest singular value together.
7 Conclusion and Discussion
@@
Limitation and future work. Due to resource constraints, our research reached the scale of ViT-L. Exploring validation on the larger models such as ViT-G or ViT-H where the assumptions behind our theory become more realistic would be necessary. Our scope of validation was also limited to CLIP-like VLMs, but Theorem 3.1 is not specific to VLMs, and investigating the applicability to other types of models such as language models would be an exciting future work direction.
Impact statement. Our method enhances the foundation models’ reliability in multiple dimensions – accuracy and confidence calibration, and many downstream applications for society can enjoy benefits from the improved reliability. However, inherent biases learned from ID data can not be removed with our method, and thus may have risk raising potential harms in real-world applications.
Acknowledgement
We thank the NeurIPS24 reviewers, Soojeong Lee and Jinho Kang, for their constructive feedback. Some parts of experiments are based on the NAVER Smart Machine Learning (NSML) platform [54]. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2022-00143911, AI Excellence Global Innovative Leader Education Program) and (No.RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST)), and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00457216, 2022R1A4A3033874). This work was also supported by the Air Force Research Laboratory under agreement number FA8750-19-2-0200, and by grants from the Defense Advanced Research Projects Agency (DARPA) through the GAILA program (award HR00111990063) and the AIDA program (FA8750-18-20018). The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright notice herein. The views and conclusions expressed in this document are those of the authors and should not be interpreted as representing the official policies or endorsements, either expressed or implied, of the Air Force Research Laboratory or the U.S. Government.
@@ Appendix A Experimental Detail
This section supplements §5 by providing detailed descriptions for experiments to enhance reproducibility.
@@ A.1 Details for numerical analysis on error bounds
We generate 1000-dimensional Gaussian random variables, where the variables have binary noisy labels ($15\%$ of random flip) for the ID train, ID test, and OOD test datasets. For the ID train set, the first 400 dimensions and the second 400 dimensions are correlated with labels, and the remaining 200 dimensions are zero-centered random noises. We build the OOD test set from the ID test set by shifting the mean of the first 400 dimensions and downscaling the second 400 dimensions in half. The remaining 200 dimensions are intact. For example, the feature $x_{2}$ in Figure A is perfectly correlated with labels across train and test in both ID and OOD environments, while the correlation between feature $x_{1}$ and labels vanish in OOD environment. We train the three-layer multi-layer perceptron networks with four different learning objectives adopting regularization terms for calibration ($\mathcal{L}_{\text{SD}}$) and for the smallest singular value ($\mathcal{L}_{\text{OC}}$) with varying regularization magnitudes (111 models in total): (i) without regularization (Baseline) (ii) with $\mathcal{L}_{\text{SD}}$, (iii) with $\mathcal{L}_{\text{OC}}$, and (iv) with $\mathcal{L}_{\text{SD}}$ and $\mathcal{L}_{\text{OC}}$. For $\mathcal{L}_{\text{OC}}$, we use an orthogonal constraint over the last weight matrix. For $\mathcal{L}_{\text{SD}}$, we adopt Born-Again-Network (BAN)-style self-distillation [13]. After training, we measured $\sigma_{min}$ and ECE on the ID test set and measured the mean squared error (MSE) and ECE on OOD test set. The visualization results of Figure 3 show the correlation between the OOD MSE (or OOD ECE) and its upper bound (average of the standardized444computed by subtracting the mean and dividing with standard deviation of all 111 models’ metric values smallest singular value of ID feature representation and standardized ID ECE).
@@ Appendix B Additional Evaluation Results
In addition to the comparisons of our CaRot with zero-shot, naive fine-tuning, and robust fine-tuning methods in §5, we provide results when applying post-hoc techniques for robustness (weight ensembling) and calibration (temperature scaling). Moreover, we compare our self-distillation-based soft label with uniform constant label smoothing.
@@ B.1 Comparing approaches
B.2 Additional result and discussion
Results with WiSE are reported in Table B and C. We observe that weight ensembling, which aims to align zero-shot models and fine-tuned models in the model weight space, boosts the overall performance. Still, CaRot shows superior results to the baselines.
Results with TS can be found in Table D. Aligning with previous observations, applying TS significantly improves ID ECE. However, since TS adjusts the logits assuming that all data instances would have a similar extent of overconfidence (or underconfidence) issue, its positive effect is limited under distribution shifts. Hence, TS does not guarantee to attain low calibration error on OOD datasets. It is noteworthy that applying TS even harms OOD ECE of CaRot (compare results with Table 3), which is already calibrated using data-dependent confidence adjustment during fine-tuning.
Results comparing with LS are reported in Table E. In accordance with its capability of improving calibration, LS successfully reduces calibration errors. We observe that our adopted approach, EMA SD, remarkably outperforms LS, especially on OOD ECE as well as in ID ECE. While LS addresses the over-confidence problem during train time by dispersing the concentrated confidence on the target label to non-target labels with a constant amount, self-distillation shows a similar behavior but in a dynamic approach, considering the diversity of data instances [71]. We interpret that reflecting the difficulty of input batch and distilling such information is crucial to achieving robust calibration.
@@ B.2 Additional result and discussion
Other types of regularization for larger smallest singular value. We adopted an orthogonality constraint over visual projection matrix $W_{v}\in\mathbb{R}^{d_{v}\times r}$ via $||W_{v}^{T}W_{v}-\mathbf{I}||_{F}$ term which pursues the larger effective rank of the visual projection matrix. While we adopt this kind of indirect soft constraint to achieve balanced performance over both ID and OOD generalization, one may wonder about 1) the implication of increased effective rank of $W_{v}$ on the smallest singular value of the covariance matrix and 2) the possibility of leveraging more direct constraints to increase the smallest singular value of the input covariance matrix. This paragraph answers those questions.
Given $n$ samples of $d$-dimensional visual features and $r$ as a pre-defined projection dimension, let $\tilde{Z}_{v}\in\mathbb{R}^{n\times r}$ and $\hat{Z}_{v}\in\mathbb{R}^{n\times r}$ denote visual representations obtained by an arbitrary projection matrix $W_{v}\in\mathbb{R}^{d_{v}\times r}$ and an orthogonal projection matrix $O_{v}\in\mathbb{R}^{d_{v}\times r}$ multiplied with pre-projected visual feature $Z_{v}\in\mathbb{R}^{n\times d_{v}}$, respectively. We can show that the rank of the covariance matrix from $\hat{Z}_{v}$ is always greater or equal to that of $\tilde{Z}_{v}$ as below,
$\text{rank}(\tilde{Z}_{v}^{T}\cdot\tilde{Z}_{v})$ $=\text{rank}(\tilde{Z}_{v})$ (6)
$=\text{rank}(Z_{v}\cdot W_{v})$ (7)
$\leq\text{rank}(Z_{v}\cdot O_{v})$ (8)
$=\text{rank}(\hat{Z}_{v})$ (9)
$=\text{rank}(\hat{Z}^{T}_{v}\cdot\hat{Z}_{v})$ (10)
During minimizing $||W_{v}^{T}W_{v}-\mathbf{I}||_{F}$, the visual projection matrix $W_{v}$ becomes closer to an orthogonal matrix $O_{v}$, and the effective rank of the visual representations’ covariance matrix increases. A larger effective rank implies a non-diminishing (relatively larger) smallest singular value of a matrix, which justifies our implementation of the method.
Meanwhile, we can adopt other constraint terms alternative to orthogonality constraint over visual projection matrix. Using the negative value of the smallest singular value of the visual representation’s covariance matrix as an additional loss term may be the most natural candidate. The result is provided in Table A. While both SVD and orthogonality constraint methods show remarkably better performance in terms of OOD classification and calibration, SVD requires much heavier computation compared with the orthogonality constraint term. For simplicity, we advocate the use of orthogonality term thus. Exploring other types of efficient constraint terms will be a promising future work direction.
Table A: Comparison between SVD-based regularization and the orthogonality constraint term. Both terms are effective in terms of OOD generalization and calibration, but SVD requires a much heavier computation.
ID OOD ID-OOD Gap
Time Complexity Acc. ($\uparrow$) ECE ($\downarrow$) Acc. ($\uparrow$) ECE ($\downarrow$) Acc. ($\downarrow$) ECE ($\downarrow$)
FLYP - 82.69 0.0635 59.40 0.1836 23.29 0.1201
CaRot (SVD) $\mathcal{O}(D^{2}N+D^{3})$ 83.05 0.0536 62.40 0.0770 20.65 0.0234
CaRot (Ours) $\mathcal{O}(D^{2}N+D^{2})$ 83.13 0.0470 62.55 0.0791 20.58 0.0321
Table B: Accuracy on ImageNet and distribution shifts using WiSE-FT [61]. We select the optimal ensembling coefficient (i.e., $\alpha$) for each method.
@@
Table C: ImageNet ECE results with WiSE-FT [61]. Along with Table B, we report the ECE results.
@@
Table D: ImageNet ECE results with temperature scaling (TS). Supplement to Table 2 and 3, we provide results applying TS. Note that TS is a post-hoc method and does not affect accuracy. The temperature is selected using IN ECE for each method.
@@ B.3 Detailed results from main paper
Coefficient terms ablation. In Table G and F, we present the detailed ablation results of hyperparameters associated with the methodologies addressed in our paper. These results supplement Table 5.
Different VLM backbones. We provide the full results of Table 7 in Table H and I.
@@
Table E: Comparison on LS and EMA SD. We compare the impact of LS and EMA SD with $\mathcal{L}_{\text{MCL-con}}$ as calibration regularization.
@@
Table F: Orthogonality constraint hyperparameter. We report the impact of the orthogonality constraint term of the CaRot objective by ablating its strength coefficient $\lambda_{\text{OC}}$. We set our final value as 0.2, tuning based on ID ECE.
@@
Table G: Self-distillation term hyperparameter. We report the impact of the EMA SD term of CaRot objective by ablating its strength coefficient $\lambda_{\text{SD}}$. We set our final value as 1.5, tuning based on $\sigma_{min}$.
@@
Table H: ImageNet results on CLIP ResNet50
@@
Table I: ImageNet results on CLIP ViT-L/14
@@
Appendix C Proof and further discussion
In this section, we provide proof for Theorem 3.1 and some discussions.
@@ Appendix C Proof and further discussion
Let $h:\mathcal{X}\rightarrow[0,1]$ be a real-valued function of structure $h(x)=\sum_{i=1}^{d}h_{i}(x[i])$ where $h_{i}$ is an arbitrary one-dimensional function, and $h$ is in a hypothesis class $\mathcal{H}$ that has pseudo dimension $\mathcal{P}dim(\mathcal{H})=d_{h}$, $\hat{\mathcal{D}}_{\text{ID}}$ be an $N$-size empirical distribution on ID domain. If $(x[1],...,x[d])$ have matching marginals for ID and OOD, and $(x[i],x[j])$ is a bi-variate Gaussian for every $i,j\in[d]$, then for any $\delta\in(0,1)$ and for all $h$, the following bounds hold with probability at least $1-\delta$:
@@
where $\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}}{:=}\mathbb{E}_{\mathcal{D}_{\text{ID}}}[\tilde{x}\tilde{x}^{T}]$ is a covariance matrix with a strictly positive minimum singular value of $d$-dimensional normalized input $\tilde{x}=(\tilde{x}[1],...,\tilde{x}[d])$, where $\tilde{x}[i]{:=}(x[i]-\mathbb{E}[x[i]])\text{Var}(x[i])^{-1/2}$ and $\sigma_{\text{min}}(M)$ is the smallest singular value of a matrix $M\in\mathbb{R}^{d_{1}\times d_{2}}$.
@@
$\varepsilon_{\mathcal{D}_{\text{OOD}}}(h)$ $\leq\varepsilon_{\mathcal{D}_{\text{OOD}}}(h^{*})+\varepsilon_{\mathcal{D}_{\text{OOD}}}(h,h^{*})$
$=\varepsilon_{\mathcal{D}_{\text{OOD}}}(h^{*})+\varepsilon_{\mathcal{D}_{\text{OOD}}}(h,h^{*})-\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h^{*})+\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h^{*})$
$\leq\varepsilon_{\mathcal{D}_{\text{OOD}}}(h^{*})+|\varepsilon_{\mathcal{D}_{\text{OOD}}}(h,h^{*})-\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h^{*})|+\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h^{*})$
$\leq\varepsilon_{\mathcal{D}_{\text{OOD}}}(h^{*})+\varepsilon_{\mathcal{D}_{\text{ID}}}(h,h^{*})+sd_{\mathcal{H}}(\mathcal{D}_{\text{OOD}},\mathcal{D}_{\text{ID}})$
$\leq\varepsilon_{\mathcal{D}_{\text{OOD}}}(h^{*})+\varepsilon_{\mathcal{D}_{\text{ID}}}(h)+\varepsilon_{\mathcal{D}_{\text{ID}}}(h^{*})+sd_{\mathcal{H}}(\mathcal{D}_{\text{OOD}},\mathcal{D}_{\text{ID}})$
@@
where $\tilde{\Sigma}_{\mathcal{D}_{\text{ID}}}:=\mathbb{E}_{\mathcal{D}_{\text{ID}}}[\tilde{x}\tilde{x}^{T}]$ is a covariance matrix of the $d$-dimensional nomarlized input $\tilde{x}=(\tilde{x}[1],...,\tilde{x}[d])$ where $\tilde{x}[i]:=(x[i]-\mathbb{E}[x[i]])\text{Var}(x[i])^{-1/2}$ and $\sigma_{\text{min}}(M)$ is the minimum singular value of a matrix $M\in\mathbb{R}^{d_{1}\times d_{2}}$.